OpenStudy (aundiewear):

Consider the following. u = (5, −4), v = (6, 4) (a) Find u · v. (b) Find the angle between u and v to the nearest degree. θ = for a I got 14 which is right but I don't get how to get the θ ??

OpenStudy (trojanpoem):

u . v = (5,-4).(6,4) = 30 - 16 = 14 u.v = |u||v|costheta costheta = 14/|u|*|v| |u| = sqrt(5^2 + 4^2) = |v| = sqrt(6^2 + 4^2) =

OpenStudy (aundiewear):

so would it be 37?

OpenStudy (trojanpoem):

72.35 degrees

OpenStudy (aundiewear):

wait how??

OpenStudy (aundiewear):

it would be costheta = 7/1066

OpenStudy (aundiewear):

ohh i got it thanks a bunch

OpenStudy (trojanpoem):

Any time.